What is the annual return for 2015 on the following four investments?
13. Stock A
Stock price per share January 1, 2015: $22.00
Annual Dividend: $ 1.10
Stock price per share December 31, 2015: $30.00

_____________________________________________________________________
14. Bond A
Bond price January 1, 2015: $ 975.00
Annual coupon payment: $ 70.00
Bond price December 31, 2015: $1,050.00

_____________________________________________________________________

_____________________________________________________________________
15. Stock B
Stock price per share January 1, 2015: $72.00
Annual Dividend: $ 2.55
Stock price per share December 31, 2015: $77.00
_____________________________________________________________________

_____________________________________________________________________
16. Bond B
Bond price January 1, 2015: $1,000.00
Annual coupon payment: $ 80.00
Bond price December 31, 2015: $1,200.00
Answer #1

13. Annual Return of Stock A =(Stock Price after 1 year -Stock Price now +Dividend)/Stock Price now =(30-22+1.10)/22=41.36%

14. Bond A Return =(Bond Price after 1 year +Annual Coupon-Bond Price now)/Bond Price Now =(1050+70-975)/975 =14.87%

15. Annual Return of Stock A =(Stock Price after 1 year -Stock Price now +Dividend)/Stock Price now =(77-72+2.55)/72=10.49%

16. Bond A Return =(Bond Price after 1 year +Annual Coupon-Bond Price now)/Bond Price Now
=(1200+80-1000)/1000 =28.00%